OpenStudy (anonymous):

find the value of the 3x3 matrix [-6,-3,-5] [ 5, 7, -6] [3, 9, 9]

OpenStudy (anonymous):

find the inverse of the matrix if it exists [3, -1] [-10 ,9]

OpenStudy (anonymous):

solve the system by using a matrix equation 2x-3y=3 5x-7y=9

OpenStudy (anonymous):

@amistre64 @help123please. @Ashleyisakitty

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

can you help with these?

terenzreignz (terenzreignz):

Value of a 3x3 matrix?

OpenStudy (anonymous):

yea the first one

terenzreignz (terenzreignz):

You mean determinant?

OpenStudy (anonymous):

um im not sure yea probably

OpenStudy (anonymous):

my choices are -635 -631 -633 and -636

OpenStudy (anonymous):

i did the work and i think its -635 but needed a second opinion

terenzreignz (terenzreignz):

It isn't. Could you show me how you did it?

OpenStudy (anonymous):

i multiplied the numbers diagnoly

OpenStudy (anonymous):

then added

terenzreignz (terenzreignz):

Yeah.. so show me. Need to see where you made the error

OpenStudy (anonymous):

ok [-6,-3,-5] [ 5, 7, -6] [3, 9, 9] so it would be -6* 9*-6 +-5*7*3+ 5*-3*9 right?

terenzreignz (terenzreignz):

That results in 84...

OpenStudy (anonymous):

yea im noy sure.. is that how you do it though?

terenzreignz (terenzreignz):

No... think of it this way... You have a matrix... \[\Large \left[\begin{matrix}a & b & c\\d&e&f\\g&h&i\end{matrix} \right]\]

OpenStudy (anonymous):

okay

terenzreignz (terenzreignz):

To get its determinant... copy the first and second columns and attach them to the right... \[\Large \left[\begin{matrix}a & b & c\\d&e&f\\g&h&i\end{matrix} \right|\left.\begin{matrix}a&b\\d&e\\g&h\end{matrix}\right]\]

terenzreignz (terenzreignz):

Catch me so far?

OpenStudy (anonymous):

yes i think so

terenzreignz (terenzreignz):

Okay, now you multiply diagonally, and add the products... \[\Large \left[\begin{matrix}\color{red}a & \color{green}b & \color{blue}c\\d&\color{red}e&\color{green}f\\g&h&\color{red}i\end{matrix} \right|\left.\begin{matrix} a&b\\\color{blue}d&e\\\color{green}g&\color{blue}h\end{matrix}\right]\]

OpenStudy (anonymous):

oh okay

terenzreignz (terenzreignz):

So you get... \[\Large \color{red}{aei}+ \color{green}{bfg}+\color{blue}{cdh}\] (this is not the entire determinant yet)

OpenStudy (anonymous):

i got -549

terenzreignz (terenzreignz):

And then, multiply THESE diagonals (this time, starting from the bottom) \[\Large \left[\begin{matrix}a & b & \color{red}c\\d&\color{red}e&\color{green}f\\\color{red}g&\color{green}h&\color{blue}i\end{matrix} \right|\left.\begin{matrix}\color{green}a&\color{blue}b\\\color{blue}d&e\\g&h\end{matrix}\right]\] but you NEGATE them... \[\Large -\color{red}{gec}-\color{green}{hfa}-\color{blue}{idb}\]

OpenStudy (anonymous):

okay one min

OpenStudy (anonymous):

84?

terenzreignz (terenzreignz):

don't forget the negative signs...

OpenStudy (anonymous):

so -84

terenzreignz (terenzreignz):

So that the total determinant is \[\Large \color{red}{aei}+ \color{green}{bfg}+\color{blue}{cdh} -\color{red}{gec}-\color{green}{hfa}-\color{blue}{idb}\]

terenzreignz (terenzreignz):

So, what's your answer?

OpenStudy (anonymous):

would it be -633 then?

terenzreignz (terenzreignz):

That is correct :)

OpenStudy (anonymous):

thank you! can you help with the next two then ill not bother you?

terenzreignz (terenzreignz):

Can't. I have to go now :) Sorry... I'm sure there's someone here who'd be more than willing to help :)

OpenStudy (anonymous):

okay thank you soo much!